# Emanuel Castelo Aix-Marseille Université, CNRS, LIS, 13009 Marseille, France Oscar Defrain Aix-Marseille Université, CNRS, LIS, 13009 Marseille, France Guilherme C. M. Gomes LIRMM, Université de Montpellier, CNRS, Montpellier, France Universidade Federal de Minas Gerais, Belo Horizonte, Brazil

### v2

@@ 4 Minimal extension for (total) dominating sets
First note that since the edges of $\Delta_{i+1}$ belong to $\mathcal{H}_{i+1}$ and are incident to $v_{i+1}$, they may be equal to $N[v_{i+1}]$ if $N[v_{i+1}]\subseteq V_{i+1}$, or to $N[u]$ for some neighbor $u$ of $v$ if $N[u]\subseteq V_{i+1}$. From now on, let $B$ be the set of vertices $u\in N(v_{i+1})$ such that $N[u]\in\Delta_{i+1}$, and let
$R:=\left(\bigcup\Delta_{i+1}\right)\setminus B$
be the remaining vertices that lie in some edge of $\Delta_{i+1}$. Consequently $B\subseteq N(v_{i+1})$ and $R\subseteq N(v_{i+1})\cup N^{2}(v_{i+1})$. Then, by Proposition 4, the graph $H=G_{i}[R\cup B]$ is bipartite chain. We denote by $(X,Y)$ its bipartition where $B\subseteq X\subseteq N(v_{i+1})$, and by $x_{1},\dots,x_{p}$ and $y_{1},\dots,y_{q}$ the respective orderings of $X$ and $Y$ satisfying $N(x_{j})\cap Y\subseteq N(x_{\ell})\cap Y$ and $N(y_{\ell})\cap X\subseteq N(y_{j})\cap X$ for all $j<\ell$; see Section 2 and Figure 1 for an illustration of the situation.
@@
First, note that if $x$ belongs to $T\cap R\cap N(v_{i+1})$ then it is incident to at most one edge in $\Delta_{i+1}$ (as $H$ is bipartite), namely $N[v_{i+1}]$, in which case $N[v_{i+1}]$ is a private edge of $x$. Hence, no two distinct vertices $x,y$ may lie in $T\cap R\cap N(v_{i+1})$ by minimality of $T$, yielding the first inequality.
Now, recall that $N^{2}(v_{i+1})\subseteq Y$ hence that one of $N(y)\cap X\subseteq N(y^{\prime})\cap X$ or $N(y)\cap X\supseteq N(y^{\prime})\cap X$ holds for any pair $y,y^{\prime}\in N^{2}(v_{i+1})$. Since edges of $\Delta_{i+1}$ are of the form $N[x]$ for some $x\in X$, no two distinct $y,y^{\prime}$ can have private edges in $\Delta_{i+1}$ simultaneously. Hence, $|T\cap N^{2}(v_{i+1})|\leq 1$ as claimed. ∎
@@
Let $T$ be a minimal transversal of $\Delta_{i+1}$. Then exactly one of the following holds:
@@
    $T\subseteq B$ in which case $T=B$;
@@
    $T\subseteq R$ in which case $|T|\leq 2$; or
@@
First, note that all the cases are mutually distinct. Hence, we may consider some $T\in Tr(\Delta_{i+1})$, assume that $T\neq\{v_{i+1}\}$, and prove the conclusion for each of Items 24 Note that each $x\in X$ hits a single edge $N[x]$ in $\Delta_{i+1}$, as $X$ is an independent set. Hence, if $T\subseteq B$ then $T=B$ concluding Item 2. If $T\subseteq R$ then we have that $|T|\leq 2$ by Proposition 9 deriving Item 3. We are thus left with proving Item 4. Suppose that $T\cap B\neq\emptyset$ and $T\cap R\neq\emptyset$. Note that if $x\in X\cap R$, then it may be incident to at most one edge in $\Delta_{i+1}$, namely $N[v_{i+1}]$. However, since $T\cap B\neq\emptyset$ by assumption, and since each element in $B$ is adjacent to $v_{i+1}$, the edge $N[v_{i+1}]$ may not be the private edge of such an $x$ in $T$. Hence $T\cap X\cap R=\emptyset$, that is, $T\cap R\subseteq Y$. By Proposition 9 there is exactly one vertex, call it $y$, belonging to $T\cap Y$. Finally, note that in addition to $N[v_{i+1}]$, each $b\in B$ hits a single edge $N[b]\in\Delta_{i+1}$ that only $b$ can hit in $X$. Hence $B\setminus N(r)\subseteq T$ and Item 4 follows. ∎
@@ 5 Minimal connected dominating sets
If $G$ is chordal bipartite then $\mathcal{S}(G)$ has conformality $5$.
@@
Let us put $\mathcal{H}:=\mathcal{S}(G)$. Assume that $\mathcal{H}$ is not of conformality 55{5}5, i.e., there is a subset $X\subseteq V(\mathcal{H})$ that is not contained in a hyperedge of $\mathcal{H}$, but each of its subsets $Y\subseteq X$ of size at most $5$ are contained in a hyperedge of $\mathcal{H}$. Here, by contained we mean as a subset. We consider $X=\{x_{1},\dots,x_{p}\}$ of minimum cardinality. Then ${p\geq 6}$ as otherwise, by assumption, $X$ would be contained in a hyperedge of $\mathcal{H}$. Moreover, for every $x_{i}\in X$, note that there must exist a hyperedge $E_{i}$ of $\mathcal{H}$ such that $E_{i}\cap X=X\setminus\{x_{i}\}$: we prove this depending on the value of $p$. In the first case, if $p\geq 7$, the nonexistence of such $E_{i}$ implies that $X^{\prime}=X\setminus\{x_{i}\}$ is not contained in a hyperedge of $\mathcal{H}$, and still every subset $Y\subseteq X^{\prime}$ of size at most $5$ is, contradicting the minimality of $X$; if, however, $p=6$ by definition we have that every $Y\subset X$ of size 5 is contained in some $E_{Y}\in E(\mathcal{H})$, but $X\nsubseteq E_{Y}$ since $\mathcal{H}$ is not of conformality 5.
Since $X$ has cardinality at least $6$, there exist at least three vertices $a_{1},a_{2},a_{3}$ in $X$ lying in one side of the bipartition of $G$, call it $A$. Let $E_{1}$ be an edge of $\mathcal{H}$, i.e., a minimal separator of $G$ that does not contain $a_{1}$, which exists by our previous arguments. By the minimality of $E_{1}$ and Proposition 7 there are at least two components that are close to $E_{1}$. Let $C_{1}$ be one that does not contain $a_{1}$. By Proposition 8 there exists $c_{1}$ in $C_{1}$ that is adjacent to both $a_{2}$ and $a_{3}$. However note that $c_{1}$ is not adjacent to $a_{1}$ as we chose $C_{1}$ to be the component not containing $a_{1}$.
@@ 6 Limitations of the sequential method
Figure 2: The construction of $H$, where vertices are ordered arbitrarily except for $u^{\star}$ and $v^{\star}$ which are put last, in this order. Color classes $V_{1},\dots,V_{k}$ induce independent sets in $H$. The dashed edge denote an edge of $G$ that is not part of $H$. Black vertices represent $T^{\star}$. Note that there is one copy of the depicted gadgets for each color class, vertex, and edge of $G$.
@@
For every edge $uv$ in $G$ we construct a subdivided claw $y_{1},y_{2},y_{3},y_{4},u,v$, with $N(y_{2})=\{y_{1},y_{3}\}$, $N(y_{4})=\{y_{3},u\}$, $N(y_{5})=\{y_{3},v\}$ and no other edges touch the $y$ vertices. Also, add $y_{2}$ and $y_{3}$ to $T^{\star}$. The objective of this gadget is to forbid the selection of both endpoints $u$ and $v$ of an edge in a minimal dominating set containing $T^{\star}$. Note that we construct one such gadget per edge of $G$.
@@
Finally, we create two additional vertices $u^{\star}$ and $v^{\star}$ with $u^{\star}$ adjacent to each $c_{i}$, $1\leq i\leq k$, and $v^{\star}$ only adjacent to $u^{\star}$. We fix the ordering of vertices of $H$ to be arbitrary but ending with $u^{\star},v^{\star}$ in this order.
This completes the construction of $H$. It has $O(n+m)$ vertices and edges. Moreover, it is easily seen that $H$ is bipartite, by taking $V(G)$ on one side of the bipartition, and noting that except for $u^{\star}$ and $v^{\star}$ the gadgets only create even paths between pairs of vertices attached to $V(G)$. Then $u^{\star}$ is made adjacent to the $c_{j}$’s which lie in the opposite partition of $V(G)$, and $v^{\star}$ is a pendant vertex. Thus they do not create odd cycles. See also Figure 2 for a representation of $H$ in layers where each edge is between consecutive layers.
Let us define $\mathcal{H}:=\mathcal{N}(H)$ and show that $T^{\star}$ is a minimal transversal of $\mathcal{H}_{n-2}=\mathcal{H}\left[V(H)\setminus\{u^{\star},v^{\star}\}\right]$. First note that all neighborhoods except for those of $c_{i}$’s, $u^{\star}$ and $v^{\star}$ are included in the $n-2$ first vertices, hence belong to $\mathcal{H}_{n-2}$. Moreover, all these neighborhoods are intersected by $T^{\star}$ as every gadget is dominated by $T^{\star}$ by construction, and each vertex $v\in V(G)$ is dominated by the vertex $z_{3}$ of its corresponding vertex gadget. Furthermore, $T^{\star}$ is minimal with that property as each of its element $t$ has a private neighbor in its corresponding gadget, whose edge is thus only intersected by $t$ in $T^{\star}$.
Let us now consider the different possibilities for a minimal transversal $T$ of $\mathcal{H}_{n-1}$ to contain $T^{\star}$ as a subset. Note that only the neighborhoods of $c_{i}$ for each $1\leq i\leq k$ are to be dominated. To do so, we either pick $u^{\star}$, or must select a subset $X$ of $V(G)\cup C$ in order to minimally hit $N[c_{i}]$ for all $1\leq i\leq k$. Note however that as $x_{4}$ is the only private neighbor of $x_{3}$ in the color class gadget of $V_{i}$, $X$ cannot contain $c_{i}$ nor $x_{1}$ for any color. Hence $X\subseteq V(G)$. Note also that no two distinct elements of $X$ intersect a same color class $V_{i}$ as their only private neighbor would be $c_{i}$, and hence one of the two would be redundant. Moreover by the edge gadget, since $y_{4}$ and $y_{5}$ are the only two private neighbors of $y_{3}$, $X$ cannot contain both $u$ and $v$ for any edge $uv$ of $G$. Thus $X$ intersects each color class precisely once, and does not contain two vertices $u,v$ with $uv\in E(G)$. We conclude that $X$ is a multicolored independent set of $G$, hence that $T^{\star}$ can be extended into a minimal transversal of $\mathcal{H}_{n-1}$ using a different set than $\{u^{*}\}$ if and only if $G$ contains a multicolored independent set.
@@
$\Delta_{n-1}:=\{E\in\mathcal{H}_{n-1}:E\cap T^{\star}=\emptyset\}=\{N[c_{i}]:1\leq i\leq k\}.$
Since $T^{\star}$ is not a transversal of $\mathcal{H}_{n-1}$ we are in the second case. Moreover, by the above discussion, each such $X\in Tr(\Delta_{n-1})$ is either $\{u^{*}\}$ or is a multicolored independent set of $G$. We thus need to show that any multicolored independent set of $G$ yields a children to conclude the proof.
Let $X$ be a multicolored independent set of $G$ and suppose that $T^{\star}\cup X$ is not a child of $(T^{*},n-2)$. Then during the computation of the parent of $T^{\star}\cup X$, some element $v\in T^{*}$ is removed. Hence such a $v$ belongs to one of the edge, vertex, of color class gadgets. But for these element to lose a private neighbor it should be that $X$ contains two elements of a same color class, or two endpoints of an edge, a contradiction. ∎
@@
7 Trans-Enum and bipartite graphs

### v3

@@ 4 Minimal extension for (total) dominating sets
First note that since the edges of $\Delta_{i+1}$ belong to $\mathcal{H}_{i+1}$ and are incident to $v_{i+1}$, they may be equal to $N[v_{i+1}]$ if $N[v_{i+1}]\subseteq V_{i+1}$, or to $N[u]$ for some neighbor $u$ of $v_{i+1}$ if $N[u]\subseteq V_{i+1}$. Note that in the later case, the set $N[u]$ may contain vertices at distance 2 from $v_{i+1}$. From now on, let $B$ be the set of vertices $u\in N(v_{i+1})$ such that $N[u]\in\Delta_{i+1}$, and let
$R\coloneq\left(\,\bigcup\{E:E\in\Delta_{i+1}\}\right)\setminus B$
[…]
@@
First, note that if $x$ belongs to $Z\cap R\cap N(v_{i+1})$ then it is incident to at most one edge in $\Delta_{i+1}$ (as $H$ is bipartite), namely $N[v_{i+1}]$, in which case $N[v_{i+1}]$ is a private edge of $x$. Hence, no two distinct vertices $x,y$ may lie in $Z\cap R\cap N(v_{i+1})$ by minimality of $Z$, yielding the first inequality.
Now, recall that $N^{2}(v_{i+1})\subseteq Y$ hence that one of $N(y)\cap X\subseteq N(y^{\prime})\cap X$ or $N(y)\cap X\supseteq N(y^{\prime})\cap X$ holds for any pair $y,y^{\prime}\in N^{2}(v_{i+1})$. Since edges of $\Delta_{i+1}$ are of the form $N[x]$ for some $x\in X$, no two distinct $y,y^{\prime}$ can have private edges in $\Delta_{i+1}$ simultaneously. Hence, $|Z\cap N^{2}(v_{i+1})|\leq 1$ as claimed. ∎
@@
Let $Z\in Tr(\Delta_{i+1})$. Then exactly one of the following holds:
@@
    $Z\subseteq B$ in which case $Z=B$;
@@
    $Z\subseteq R$ in which case $|Z|\leq 2$; or
@@
First, note that all the cases are mutually distinct. Hence, we may consider some $Z\in Tr(\Delta_{i+1})$, assume that $Z\neq\{v_{i+1}\}$, and prove the conclusion for each of Items 24 Note that each $x\in X$ hits a single edge $N[x]$ in $\Delta_{i+1}$, as $X$ is an independent set. Hence, if $Z\subseteq B$ then $Z=B$ concluding Item 2. If $Z\subseteq R$ then we have that $|Z|\leq 2$ by Proposition 9 deriving Item 3. We are thus left with proving Item 4. Suppose that $Z\cap B\neq\emptyset$ and $Z\cap R\neq\emptyset$. Note that if $x\in X\cap R$, then it may be incident to at most one edge in $\Delta_{i+1}$, namely $N[v_{i+1}]$. However, since $Z\cap B\neq\emptyset$ by assumption of this last case, and since each element in $B$ is adjacent to $v_{i+1}$, the edge $N[v_{i+1}]$ may not be the private edge of such an $x$ in $Z$. Hence $Z\cap X\cap R=\emptyset$, that is, $Z\cap R\subseteq Y$. By Proposition 9 there is exactly one vertex, call it $y$, belonging to $Z\cap Y$. Finally, note that in addition to $N[v_{i+1}]$, each $b\in B$ hits a single edge $N[b]\in\Delta_{i+1}$ that only $b$ can hit in $X$. Hence $B\setminus N(r)\subseteq Z$ and Item 4 follows. ∎
@@ 5 Minimal connected dominating sets
If $G$ is chordal bipartite then $\mathcal{S}(G)$ has conformality 555.
@@
Let us put $\mathcal{H}\coloneq\mathcal{S}(G)$. Assume that $\mathcal{H}$ is not of conformality 5{5}5, i.e., there is a subset $X\subseteq V(\mathcal{H})$ that is not contained in a hyperedge of $\mathcal{H}$, but each of its subsets $Y\subseteq X$ of size at most 555 are contained in a hyperedge of $\mathcal{H}$. Here, by contained we mean as a subset. We consider $X=\{x_{1},\dots,x_{p}\}$ of minimum cardinality. Then ${p\geq 6}$ as otherwise, by assumption, $X$ would be contained in a hyperedge of $\mathcal{H}$. Moreover, for every $x_{i}\in X$, note that there must exist a hyperedge $E_{i}$ of $\mathcal{H}$ such that $E_{i}\cap X=X\setminus\{x_{i}\}$: we prove this depending on the value of $p$. In the first case, if $p\geq 7$, the nonexistence of such $E_{i}$ implies that $X^{\prime}=X\setminus\{x_{i}\}$ is not contained in a hyperedge of $\mathcal{H}$, and still every subset $Y\subseteq X^{\prime}$ of size at most 555 is, contradicting the minimality of $X$; if, however, $p=6$, by definition we have that every $Y\subsetneq X$ of size 5 is contained in some $E_{Y}\in E(\mathcal{H})$, but $X\nsubseteq E_{Y}$ since $\mathcal{H}$ is not of conformality 5.
Since $X$ has cardinality at least 666, there exist at least three vertices $a_{1},a_{2},a_{3}$ in $X$ lying in one side of the bipartition of $G$, call it $A$. Let $E_{1}$ be an edge of $\mathcal{H}$, i.e., a minimal separator of $G$ that does not contain $a_{1}$, which exists by our previous arguments; moreover, $E_{1}\cap X=X\setminus\{a_{1}\}$, implying $a_{2},a_{3}\in E_{1}$. By the minimality of $E_{1}$ and Proposition 7 there are at least two components that are close to $E_{1}$. Let $C_{1}$ be one that does not contain $a_{1}$. By Proposition 8 there exists $c_{1}$ in $C_{1}$ that satisfies $N(c_{1})\cap A=E_{1}\cap A$, and so $c_{1}$ is adjacent to both $a_{2}$ and $a_{3}$. However note that $c_{1}$ is not adjacent to $a_{1}$ as we chose $C_{1}$ to be the component not containing $a_{1}$.
@@ 6 Limitations of the sequential method
Figure 2: The construction of $H$, where vertices are ordered arbitrarily except for $\alpha$ and $\beta$ which are put last, in this order. Color classes $V_{1},\dots,V_{k}$ induce independent sets in $H$. The dashed edge denote an edge of $G$ that is not part of $H$. Black vertices represent $T^{\star}$. Note that there is one copy of the depicted gadgets for each color class, vertex, and edge of $G$.
@@
For every edge $uv$ in $G$ we construct a subdivided claw $y_{1},y_{2},y_{3},y_{4},y_{5},u,v$, with $N(y_{2})=\{y_{1},y_{3}\}$, $N(y_{4})=\{y_{3},u\}$, $N(y_{5})=\{y_{3},v\}$ and no other edges touch the $y$ vertices. Also, we add $y_{2}$ and $y_{3}$ to $T^{\star}$. The objective of this gadget is to forbid the selection of both endpoints $u$ and $v$ of an edge in a minimal dominating set containing $T^{\star}$. Note that we construct one such gadget per edge of $G$, but that the vertices $u$ and $v$ are shared by other edge gadgets.
@@
Finally, we create two additional vertices $\alpha$ and $\beta$ with $\alpha$ adjacent to each $c_{i}$, $1\leq i\leq k$, and $\beta$ only adjacent to $\alpha$. We fix the ordering of vertices of $H$ to be arbitrary but ending with $\alpha,\beta$ in this order.
This completes the construction of $H$. It has $O(n+m)$ vertices and edges. Moreover, it is easily seen that $H$ is bipartite, by taking $V(G)$ on one side of the bipartition, and noting that except for $\alpha$ and $\beta$ the gadgets only create even paths between pairs of vertices attached to $V(G)$. Then $\alpha$ is made adjacent to the $c_{j}$’s which lie in the opposite partition of $V(G)$, and $\beta$ is a pendant vertex. Thus they do not create odd cycles. See also Figure 2 for a representation of $H$ in layers where each edge is between consecutive layers.
Let us define $\mathcal{H}\coloneq\mathcal{N}(H)$ and show that $T^{\star}$ is a minimal transversal of $\mathcal{H}_{n-2}=\mathcal{H}\left[V(H)\setminus\{\alpha,\beta\}\right]$. First note that all neighborhoods except for those of $c_{i}$’s, $\alpha$ and $\beta$ are included in the $n-2$ first vertices, hence belong to $\mathcal{H}_{n-2}$. Moreover, all these neighborhoods are intersected by $T^{\star}$ as every gadget is dominated by $T^{\star}$ by construction, and each vertex $v\in V(G)$ is dominated by the vertex $z_{3}$ of its corresponding vertex gadget. Furthermore, $T^{\star}$ is minimal with that property as each of its element $t$ has a private neighbor in its corresponding gadget, whose edge is thus only intersected by $t$ in $T^{\star}$.
Let us now consider the different possibilities for a minimal transversal $T$ of $\mathcal{H}_{n-1}$ to contain $T^{\star}$ as a subset. Note that only the neighborhoods of $c_{i}$ for each $1\leq i\leq k$ are to be dominated. To do so, we either pick $\alpha$, or must select a subset $X$ of $V(G)\cup C$ in order to minimally hit $N[c_{i}]$ for all $1\leq i\leq k$. Note however that as $x_{4}$ is the only private neighbor of $x_{3}$ in the color class gadget of $V_{i}$, $X$ cannot contain $c_{i}$ nor $x_{1}$ for any color. Hence $X\subseteq V(G)$. Note also that no two distinct elements of $X$ intersect a same color class $V_{i}$ as their only private neighbor would be $c_{i}$, and hence one of the two would be redundant. Moreover by the edge gadget, since $y_{4}$ and $y_{5}$ are the only two private neighbors of $y_{3}$, $X$ cannot contain both $u$ and $v$ for any edge $uv$ of $G$. Thus $X$ intersects each color class precisely once, and does not contain two vertices $u,v$ with $uv\in E(G)$. We conclude that $X$ is a multicolored independent set of $G$, hence that $T^{\star}$ can be extended into a minimal transversal of $\mathcal{H}_{n-1}$ using a different set than $\{\alpha\}$ if and only if $G$ contains a multicolored independent set.
@@
$\Delta_{n-1}\coloneq\{E\in\mathcal{H}_{n-1}:E\cap T^{\star}=\emptyset\}=\{N[c_{i}]:1\leq i\leq k\}.$
Since $T^{\star}$ is not a transversal of $\mathcal{H}_{n-1}$ we are in the second case. Moreover, by the above discussion, each such $X\in Tr(\Delta_{n-1})$ is either $\{\alpha\}$ or is a multicolored independent set of $G$. We thus need to show that any multicolored independent set of $G$ yields a children to conclude the proof.
Let $X$ be a multicolored independent set of $G$ and suppose that $T^{\star}\cup X$ is not a child of $(T^{\star},n-2)$. Then during the computation of the parent of $T^{\star}\cup X$, some element $v\in T^{\star}$ is removed. Hence such a $v$ belongs to one of the edge, vertex, of color class gadgets. But for these element to lose a private neighbor it should be that $X$ contains two elements of a same color class, or two endpoints of an edge, a contradiction. ∎
@@
7 Trans-Enum and bipartite graphs [^2] After the submission of this work, it has come to our attention that there is a recent stronger result on the subject [KKMO24]. We nevertheless leave the proof as is since it is simple and for self-containment.